Nova Companys total overhead cost at various levels of activity are presented below: Total Machine-Overhead Month April May June July Cost Hours 49,000190,590 39,000 $165,490 59,000215, 690 69,000 240,790 Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 39,000 machine-hour level of activity is: Utilities (variable) Supervisory salaries (fixed) Maintenance (mixed) Total overhead cost $ 50,700 49,000 65,790 $165,490 Nova Companys management wants to break down the maintenance cost into its variable and fixed cost elements Required 1. Estimate how much of the $240,790 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $240,790 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.) 2. Using the high-low method, estimate a cost formula for maintenance in the form Y=a +bX. 3. Express the companys total overhead cost in the form Y-a+ bX. 4. What total overhead cost would you expect to be incurred at an activity level of 44,000 machine-hours?
Answer #1

1

Answer is $102090

Supervisor salaries =$49000

Variable cost per Machine hr =$50700/39000=$1.30 per Mh

Utilities variable =69000*$1.30 per Mh=$89700

Maintenance Cost =240790-49000-89700=$102090

2

8duoWUAoeUnx8GQWgTkgRC97i4uAjadQii28L6X3

3

$240790=$67600+69000*2.51

4

67600+44000*2.51=$178040